# Revisiting Graph Analytics Benchmark

### v1

@@ 8.2. Algorithm & Statistics Impact
Algorithm Coverage across Platforms. All experiment codes are sourced from official documentation, third-party repositories, or implemented by us to the best of our ability. Out of the total $8$ (algorithms) $\times$ $7$ (platforms) $=56$ cases, we successfully evaluate 49 cases. An algorithm’s suitability for a given computing model (Section 3.3) affects whether it is feasible and/or efficient to implement the algorithm on different platforms. Among the 7 unimplemented cases: $\mathsf{Pregel+}$’s interface lacks support for managing variables like coreness values required by $\mathsf{CD}$ across supersteps; $\mathsf{PR}$, $\mathsf{LPA}$, $\mathsf{SSSP}$, $\mathsf{WCC}$, $\mathsf{BC}$, and $\mathsf{CD}$ cannot be implemented on $\mathsf{G}$-$\mathsf{thinker}$, a subgraph-centric platform that does not support the iterative control flow needed by these algorithms. Additionally, $\mathsf{GraphX}$ struggles with $\mathsf{LPA}$, $\mathsf{CD}$, and $\mathsf{KC}$ due to the overhead for adapting the Spark RDD APIs; $\mathsf{Pregel+}$ lacks push/pull optimizations found in newer platforms like $\mathsf{Flash}$. Consequently, it faces high computation overhead with subgraph algorithms ($\mathsf{TC}$ and $\mathsf{KC}$). To manage these inefficiencies, we use 16 machines instead of one single machine for these cases (indicated by red bars in Figure 10).
@@
Iterative Algorithms ($\mathsf{PR}$ and $\mathsf{LPA}$) perform similarly on S8-Std and S8-Diam, suggesting that these algorithms are not sensitive to diameter. However, there are noticeable differences in performance on S8-Dense. For $\mathsf{PR}$, all platforms run faster on S8-Dense as expected, as $\mathsf{PR}$ distributes its workload across vertices, and the denser dataset has fewer vertices with the same number of edges. The performance of $\mathsf{LPA}$ is affected by implementation, which requires hash tables to store vertex labels. Specifically, $\mathsf{GraphX}$ experiences significant performance degradation when merging hash tables from different vertices. In contrast, other platforms maintain a local hash table and can merge received hash tables to the local one, avoiding much redundant computation.
@@
$\mathsf{Flash}$ and $\mathsf{Pregel+}$ offer standard Pregel-like APIs that support global communication, i.e., send messages to any vertices, enabling some advanced algorithms such as HashMin (Rastogi et al., 2013) and Shiloach-Vishkin (Vishkin and Shiloach, 1982; Yan et al., 2014b; Qin et al., 2014) for $\mathsf{WCC}$ and reducing iteration rounds significantly. $\mathsf{Block}$-$\mathsf{centric}$ models like $\mathsf{Grape}$ can even directly call sequential implementation, e.g., Disjoint Set, for dynamic $\mathsf{WCC}$. In contrast, some $\mathsf{vertex}$-$\mathsf{centric}$ platforms (e.g., $\mathsf{GraphX}$) and the $\mathsf{edge}$-$\mathsf{centric}$ model can only send messages to neighbors.
Unlike other sequential algorithms, the performance of $\mathsf{CD}$ exhibits varying trends across different platforms when executed on denser datasets. Since the coreness value could be very large (especially in S8-Dense) and multiple iterations may be required for each coreness value, $\mathsf{GraphX}$ is extremely slow even we use 16 machines. Additionally, for each coreness value, platforms like $\mathsf{GraphX}$ and $\mathsf{PowerGraph}$ need to activate all vertices, while $\mathsf{Flash}$ and $\mathsf{Ligra}$ can maintain the activated vertices, resulting in better performance.
@@ 8.3. Scalability Sensitivity
Varying Number of Threads. We first examine the scale-up performance on a single machine with increasing threads: $1$, $2$, $4$, $8$, $16$, and $32$. $\mathsf{GraphX}$ requires at least four threads for $\mathsf{PR}$ and two threads for $\mathsf{SSSP}$ to operate effectively, so we only exclude $\mathsf{TC}$.
Figure 11 illustrates the scale-up performance of $\mathsf{PR}$, $\mathsf{SSSP}$, and $\mathsf{TC}$ across platforms. All platforms exhibit a clear decreasing trend in execution time as more threads are used. The detailed scaling factors, calculated as the ratio of best performance to single-thread performance, are presented in Table 10. $\mathsf{Grape}$ achieves the best scale-up performance, with a scaling factor of up to 37×37$\times$37 ×, followed by $\mathsf{Ligra}$, $\mathsf{G}$-$\mathsf{thinker}$ and $\mathsf{Pregel+}$, which have an average scaling factor of 24.9×24.9$\times$24.9 ×, 23.3×23.3$\times$23.3 × and 19.6×19.6$\times$19.6 × . The scaling factors for $\mathsf{Flash}$, $\mathsf{GraphX}$ and $\mathsf{PowerGraph}$ range between 1×1$\times$1 × and 12×12$\times$12 ×.
@@
The scale-up performance varies significantly across the three algorithms. The scaling factor for $\mathsf{TC}$ is higher than $\mathsf{PR}$ and $\mathsf{SSSP}$. This confirms our previous categorization: $\mathsf{TC}$ requires no synchronization, allowing all computations to be parallelized, while $\mathsf{PR}$ consists of $k$ iterations and $k$ synchronizations. Sequential algorithms like $\mathsf{SSSP}$ require multiple synchronizations, making its scalability the worst among the three algorithms. Besides, the scale-up performance is sensitive to dataset characteristics. Most platforms scale better on S8-Dense but perform worse on S8-Diam than on S8-Std.
@@
Varying Number of Machines. We evaluate scale-out performance by increasing the number of machines to $2$, $4$, $8$, and $16$, using larger S9 datasets. $\mathsf{GraphX}$ fails to compute $\mathsf{SSSP}$ on S9-Diam, and $\mathsf{GraphX}$ and $\mathsf{PowerGraph}$ encounter Out-Of-Memory errors on $\mathsf{TC}$. $\mathsf{Ligra}$ is not tested as it runs only on a single machine. Figure 11 shows the performance, and scaling factors, calculated as the ratio of best to single-machine performance, are summarized in Table 11.
@@
All platforms show worse scale-out performance than scale-up, likely due to the high overhead of inter-machine communication. $\mathsf{Pregel+}$ performs significantly better in scale-out than the others. Notably, $\mathsf{Grape}$ performs well on a single machine, but its performance gains saturate when scaling to multiple machines due to communication overhead. The scalability trends, influenced by algorithms and dataset characteristics, are similar to those in the scale-up experiments and won’t be further discussed.
@@ 9. Platform Selection Guide
Based on the results in Section 8, Figure 14 shows the performance of all platforms across different metrics. According to the area covered by each platform, the ranking is as follows: $\mathsf{Pregel+}$ $>$ $\mathsf{Grape}$ $>$ $\mathsf{GraphX}$ $>$ $\mathsf{G}$-$\mathsf{thinker}$ $>$ $\mathsf{Flash}$ $>$ $\mathsf{PowerGraph}$ $>$ $\mathsf{Ligra}$. $\mathsf{Pregel+}$ stands out as the top-performing platform overall, performing well on all metrics without any notable weaknesses and ranking in the top three for nearly all categories. $\mathsf{Grape}$ excels in throughput, machine speed-up, thread speed-up, and stress tests, showing strong performance and scalability. However, it scores the lowest in the three usability evaluation metrics, indicating a steep learning curve and complex API. In contrast, $\mathsf{GraphX}$ achieves the highest scores in compliance, correctness, and readability, highlighting its exceptional API usability. However, it performs poorly in thread speed-up, throughput, and stress test, indicating limited performance and scalability. $\mathsf{G}$-$\mathsf{thinker}$ delivers a balanced performance in both usability and efficiency but supports fewer algorithms, leading to a lower algorithm coverage score and a mid-tier ranking. $\mathsf{Flash}$ and $\mathsf{PowerGraph}$ also shows balanced performance, but their individual rankings for each metric are not particularly outstanding, resulting in lower overall rankings. $\mathsf{Ligra}$ performs well for single-machine tasks, especially in thread speed-up, but ranks lowest overall due to its lack of support for distributed computing.
@@
$\mathsf{GraphX}$ is ideal for users of all levels, as long as performance and scalability aren’t top priorities, thanks to its excellent usability. $\mathsf{Pregel+}$, $\mathsf{Flash}$ and $\mathsf{PowerGraph}$ are also suitable for most users, especially beginners and intermediates, due to their balanced usability and good performance with large data. For those needing strong performance, $\mathsf{G}$-$\mathsf{thinker}$ and $\mathsf{Ligra}$ provide exceptional performance. However, $\mathsf{G}$-$\mathsf{thinker}$ supports fewer algorithms and is best suited for compute-intensive tasks, while $\mathsf{Ligra}$ provides robust interfaces and extensive algorithm support, excelling in single-machine tasks. Due to its lack of support for distributed computing, $\mathsf{Ligra}$ is unsuitable for large-scale or multi-machine workloads. For maximum performance and scalability, $\mathsf{Grape}$ is recommended despite its steeper learning curve, as it delivers significant efficiency once mastered.
@@ Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.

### v2

@@ 8.2. Algorithm & Statistics Impact
Algorithm Coverage across Platforms. All experiment codes are sourced from official documentation, third-party repositories, or implemented by us to the best of our ability. Out of the total $8$ (algorithms) $\times$ 77 (platforms) =56=56 cases, we successfully evaluate 49 cases. An algorithm’s suitability for a given computing model (Section 3.3) affects whether it is feasible and/or efficient to implement the algorithm on different platforms. Among the 7 unimplemented cases: $\mathsf{Pregel+}$’s interface lacks support for managing variables like coreness values required by $\mathsf{CD}$ across supersteps; $\mathsf{PR}$, $\mathsf{LPA}$, $\mathsf{SSSP}$, $\mathsf{WCC}$, $\mathsf{BC}$, and $\mathsf{CD}$ cannot be implemented on $\mathsf{G}$-$\mathsf{thinker}$, a subgraph-centric platform that does not support the iterative control flow needed by these algorithms. Additionally, $\mathsf{GraphX}$ struggles with $\mathsf{LPA}$, $\mathsf{CD}$, and $\mathsf{KC}$ due to the overhead for adapting the Spark RDD APIs; $\mathsf{Pregel+}$ lacks push/pull optimizations found in newer platforms like $\mathsf{Flash}$. Consequently, it faces high computation overhead with subgraph algorithms ($\mathsf{TC}$ and $\mathsf{KC}$). To manage these inefficiencies, we use 16 machines instead of one single machine for these cases (indicated by red bars in Figure 10).
@@
Iterative Algorithms ($\mathsf{PR}$ and $\mathsf{LPA}$) perform similarly on S8-Std and S8-Diam, suggesting that these algorithms are not sensitive to diameter. However, there are noticeable differences in performance on S8-Dense. For $\mathsf{PR}$, all platforms run faster on S8-Dense as expected, as $\mathsf{PR}$ distributes its workload across vertices, and the denser dataset has fewer vertices with the same number of edges. The performance of $\mathsf{LPA}$ is affected by implementation, which requires hash tables to store vertex labels. Specifically, $\mathsf{GraphX}$ experiences significant performance degradation when merging hash tables from different vertices. In contrast, other platforms maintain a local hash table and can merge received hash tables into the local one, avoiding much redundant computation.
(a) $\mathsf{PR}$ (Iterative)
(b) $\mathsf{LPA}$ (Iterative)
(c) $\mathsf{SSSP}$ (Sequential)
(d) $\mathsf{WCC}$ (Sequential)
(e) $\mathsf{BC}$ (Sequential)
(f) $\mathsf{CD}$ (Sequential)
(g) $\mathsf{TC}$ (Subgraph)
(h) $\mathsf{KC}$ (Subgraph)
@@
$\mathsf{Flash}$ and $\mathsf{Pregel+}$ offer standard Pregel-like APIs that support global communication, i.e., send messages to any vertices, enabling some advanced algorithms such as HashMin (rastogi2013finding) and Shiloach-Vishkin (vishkin1982log; yan2014pregel; qin2014scalable) for $\mathsf{WCC}$ and reducing iteration rounds significantly. $\mathsf{Block}$-$\mathsf{centric}$ models like $\mathsf{Grape}$ can even directly call sequential implementation, e.g., Disjoint Set, for dynamic $\mathsf{WCC}$. In contrast, some $\mathsf{vertex}$-$\mathsf{centric}$ platforms (e.g., $\mathsf{GraphX}$) and the $\mathsf{edge}$-$\mathsf{centric}$ model can only send messages to neighbors.
Unlike other sequential algorithms, the performance of $\mathsf{CD}$ exhibits varying trends across different platforms when executed on denser datasets. Since the coreness value could be very large (especially in S8-Dense) and multiple iterations may be required for each coreness value, $\mathsf{GraphX}$ is extremely slow even when using 16 machines. Additionally, for each coreness value, platforms like $\mathsf{GraphX}$ and $\mathsf{PowerGraph}$ need to activate all vertices, while $\mathsf{Flash}$ and $\mathsf{Ligra}$ can maintain the activated vertices, resulting in better performance.
@@ 8.3. Scalability Sensitivity
Varying Number of Threads. We first examine the scale-up performance on a single machine with increasing threads: $1$, 22, 44, 88, $16$, and $32$. Since $\mathsf{GraphX}$ requires at least four threads for $\mathsf{PR}$ and two threads for $\mathsf{SSSP}$ to operate effectively, we exclude only $\mathsf{TC}$.
Figure 11 illustrates the scale-up performance of $\mathsf{PR}$, $\mathsf{SSSP}$, and $\mathsf{TC}$ across platforms. All platforms exhibit a clear decreasing trend in execution time as more threads are used. The detailed scaling factors, calculated as the ratio of best performance to single-thread performance, are presented in Table 10. $\mathsf{Grape}$ achieves the best scale-up performance, with a scaling factor of up to $37\times$, followed by $\mathsf{Ligra}$, $\mathsf{G}$-$\mathsf{thinker}$ and $\mathsf{Pregel+}$, which have an average scaling factor of $24.9\times$, $23.3\times$ and $19.6\times$ . The scaling factors for $\mathsf{Flash}$, $\mathsf{GraphX}$ and $\mathsf{PowerGraph}$ range between $1\times$ and $12\times$.
@@
The scale-up performance varies significantly across the three algorithms. The scaling factor for $\mathsf{TC}$ is higher than $\mathsf{PR}$ and $\mathsf{SSSP}$. This confirms our previous categorization: $\mathsf{TC}$ requires no synchronization, allowing all computations to be parallelized, while $\mathsf{PR}$ consists of $k$ iterations and $k$ synchronizations. Sequential algorithms like $\mathsf{SSSP}$ require multiple synchronizations, making their scalability the worst among the three algorithms. Besides, the scale-up performance is sensitive to dataset characteristics. Most platforms scale better on S8-Dense but perform worse on S8-Diam than on S8-Std.
@@
Varying Number of Machines. We evaluate scale-out performance by increasing the number of machines to $2$, 44, 88, and $16$, using larger S9 datasets. $\mathsf{GraphX}$ fails to compute $\mathsf{SSSP}$ on S9-Diam, and $\mathsf{GraphX}$ and $\mathsf{PowerGraph}$ encounter Out-Of-Memory errors on $\mathsf{TC}$. $\mathsf{Ligra}$ is not tested as it runs only on a single machine. Figure 12 shows the performance, and scaling factors, calculated as the ratio of best to single-machine performance, are summarized in Table 11.
@@
All platforms show worse scale-out performance than scale-up, likely due to the high overhead of inter-machine communication. $\mathsf{Pregel+}$ performs significantly better in scale-out than the others. Notably, $\mathsf{Grape}$ performs well on a single machine, but its performance gains saturation when scaling to multiple machines due to communication overhead. The scalability trends, influenced by algorithms and dataset characteristics, are similar to those in the scale-up experiments and won’t be further discussed.
@@ 9. Platform Selection Guide
Based on the results in Section 8, Figure 14 shows the performance of all platforms across different metrics. According to the area covered by each platform, the ranking is as follows: $\mathsf{Pregel+}$ >> $\mathsf{Grape}$ >> $\mathsf{GraphX}$ >> $\mathsf{G}$-$\mathsf{thinker}$ >> $\mathsf{Flash}$ >> $\mathsf{PowerGraph}$ >> $\mathsf{Ligra}$. $\mathsf{Pregel+}$ stands out as the top-performing platform overall, performing well on all metrics without any notable weaknesses and ranking in the top three for nearly all categories. $\mathsf{Grape}$ excels in throughput, machine speed-up, thread speed-up, and stress tests, showing strong performance and scalability. However, it scores the lowest in the three usability evaluation metrics, indicating a steep learning curve and complex API. In contrast, $\mathsf{GraphX}$ achieves the highest scores in compliance, correctness, and readability, highlighting its exceptional API usability. However, it performs poorly in thread speed-up, throughput, and stress test, indicating limited performance and scalability. $\mathsf{G}$-$\mathsf{thinker}$ delivers a balanced performance in both usability and efficiency but supports fewer algorithms, leading to a lower algorithm coverage score and a mid-tier ranking. $\mathsf{Flash}$ and $\mathsf{PowerGraph}$ also show balanced performance, but their individual rankings for each metric are not particularly outstanding, resulting in lower overall rankings. $\mathsf{Ligra}$ performs well for single-machine tasks, especially in thread speed-up, but ranks lowest overall due to its lack of support for distributed computing.
@@
$\mathsf{GraphX}$ is ideal for users of all levels, as long as performance and scalability aren not top priorities, thanks to its excellent usability. $\mathsf{Pregel+}$, $\mathsf{Flash}$ and $\mathsf{PowerGraph}$ are also suitable for most users, especially beginners and intermediates, due to their balanced usability and good performance with large-scale data. For users requiring high performance, $\mathsf{G}$-$\mathsf{thinker}$ and $\mathsf{Ligra}$ are strong choices. However, $\mathsf{G}$-$\mathsf{thinker}$ supports fewer algorithms and is best suited for compute-intensive tasks, while $\mathsf{Ligra}$ provides robust interfaces and extensive algorithm support, excelling in single-machine tasks. Due to its lack of support for distributed computing, $\mathsf{Ligra}$ is unsuitable for large-scale or multi-machine workloads. For maximum performance and scalability, $\mathsf{Grape}$ is recommended despite its steeper learning curve, as it delivers significant efficiency once mastered.
@@ Acknowledgements.
We thank the 80 anonymous human experts for their valuable contributions to the usability evaluation. Long Yuan is supported by National Key R&D Program of China 2022YFF0712100, NSFC62472225. Peng Cheng is supported by NSFC 62102149. Xuemin Lin is supported by NSFC U2241211 and U20B2046.